# The 4th Deep Learning for Code Workshop

### v2

@@ 3 Agentic System
Agent-based approaches are broadly used in code generation and repair tasks. Despite a large number of autonomous and semi-autonomous coding agents, they are not widely used in formal proofs generation and are not tailored to the Rocq specifics. To address this, we have implemented a RocqStar agentic system.
@@
To allow interaction between the agent and Rocq’s system, we develop a REST API server that provides a set of tools that are useful during the execution. We apply our domain knowledge and construct these tools to bring an agent-driven proving process as close as possible to a human-driven one. Examples of allowed function calls include checking validity of proofs, retrieving the valid prefix of given proof, gathering additional information about available entities in the context, and interacting with the context via performing commands like Print ?a to identify the type of an argument or Search ?exp to search for defined terms by a pattern. Toolset is described in detail in Appendix D. Interaction with Rocq’s system is carried out through its language server, coq-lsp [7]. To conform with a commonly used Model Context Protocol (MCP) and allow seamless agent interaction with the environment through tools, we implement an MCP server that wraps the REST API server. Among the provided tools, the most important is the proof-checking tool. It not only verifies whether a proof is valid but, in case of an error, returns detailed diagnostic information: the error message, its exact location, the valid prefix preceding the error, and the remaining goals after that prefix. This functionality allows the agent to maintain awareness of the current proof state and leverage partial proof progress.
3.1 Agent Logic
@@
The input to the agent is presented as a target theorem without a proof and a file where it was declared, see box 1 of Fig.˜2. Agent’s pipeline is logically split into two main stages: planning and execution. In the planning phase, multiple language models rigorously work out the strategy for the further implementation. During execution agents follow the plan aiming to generate the correct proof.
@@
Planning Stage  We employ the idea of multi-agent debates to generate a strategy for proving the given theorem. Specifically, two LLMs engage in a discussion: one proposes and defends an initial plan (pro LLM), while the other critiques it (con LLM); see box 2 in Fig.˜2. After several debate rounds, the entire message history is passed to a judge LLM, which determines the winner and produces the final plan. Using this procedure, we generate $k$ candidate strategies. These are then evaluated by a plan scoring LLM that assigns each a numerical score (the higher, the better). Finally, the top-$l$ plans are selected and forwarded to the Execution Stage; see box 3 of Fig.˜2.
@@
Execution Stage  For each of the selected plans, we run an executor agent that follows it step by step, invoking tools from the provided tool set — proof checker, context–inspection queries, search commands, and others, as atomic actions. Through these tool calls, the agent interacts with the environment via the MCP server. In addition to this iterative execution, we employ a reflection mechanism that monitors the progress of the proof and adjusts the strategy when necessary. We track how many consecutive erroneous proof attempts occur, and once this number exceeds a predefined threshold (set to five during evaluation), a critic model is called to assess the current proof state and identify deviations from the intended strategy. After that, we retrieve theorems along with their proofs, whose top-level goals are similar to the currently remaining goal, according to the cosine similarity between their RocqStar-ranker embeddings. We prompt the LLM to explain which tactic sequences could be helpful to finish our proof. We gather the generated criticism and send it to the replanner LLM to refine the current plan along with similar proofs and their analysis. The replanner is a separate language model that revises the plan based on the critic’s feedback and the retrieved examples. The whole message history is sent back to the executor agent. During the execution of each plan, $n$ tool calls are allowed. If valid proof is not found after $n$ tool calls, we denote the plan as failed. In this case, we ask a plan failure summarizer LLM to generate a short explanation of why the strategy execution failed and what happened during it. Then this summarized explanation is sent to the new execution stage with the next selected plan. This procedure is repeated until the correct proof is found or there are no more strategies to execute.
4 Evaluation
To evaluate our approach, partially and as a whole, we use the CoqPilot benchmarking framework. We required a dataset with a large number of human-written theorems and proofs. To compare our solution to existing ones, we decided to re-use the dataset by Kozyrev et al. [15]. It is limited to 300 theorems from the IMM project [23], which was suitable for us in terms of computational and financial costs. The theorems are partitioned into three groups, corresponding to the difficulty level. The length (in tactics) of the human-written reference proof of the theorem estimates its difficulty. The sizes of each group are chosen with respect to the initial distribution of proof lengths in the project. Final group sizes and length ranges of each group could be found in Table 2. From now on, we will refer to the described dataset as the IMM-300 dataset. For smaller ablation studies we additionally prepared IMM-50, a 50-theorem subset of IMM, constructed with the same procedure. No theorems from the dataset were present in the training set of the RocqStar ranker embedding model. Moreover, the training set only contained partial theorem goals, no initial statements. Split of both datasets into groups, details, and limitations are described in Appendix A. Computational and financial resources used for experiments are described in Appendix F.
4.1 Retrieval Mechanism
We integrate our retrieval mechanism as a ranker into CoqPilot and evaluate it on the IMM-300 dataset with different models under the hood. To assess its performance, we compare our approach against two baselines: (i) an untrained embedder model (we use gte-modernbert-base, with codebert-base yielding comparable results), and (ii) a lexical similarity baseline based on the Jaccard index. In the latter, given a target theorem statement $s_{*}$ and a set of proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, it ranks the theorems in descending order of $J(s_{*},s_{i})$, where $J(s_{*},s_{i})$ is the Jaccard-similarity index and $S_{s_{i}}$ is a set of tokens inside a statement. The statement is split into tokens by whitespaces, commas, etc. Jaccard-similarity index is semantically almost the same as the BM-25 metric and produces the same numerical results. For each theorem in the dataset, we take theorems within the same file, sort them using the ranker (Jaccard, ModernBert, or RocqStar, respectively), take the $k$ most relevant ones ($k$ is equal to 7 in our experiments) and send a request to the model to generate the completion. The chosen theorems are being sent as a few-shot prompt. Generation for each theorem is requested 12 times. If the Rocq’s system accepts any of the proofs, the theorem is considered solved. The target metric in our evaluation is the ratio of solved theorems. The evaluation results are presented in Table 1. The reported values denote mean success rates, and the $\pm$ intervals correspond to the standard deviation across three independent runs.
@@ 4.1 Retrieval Mechanism
Table 1: Model performance under different ablations across all evaluation sets.
4.2 Agentic System
We evaluate our agentic system on the IMM-300 dataset, pursuing the goal to solve as many theorems as possible. For all of the parts of the planning stage, we use the Claude 3.5 Sonnet model, performing two rounds of debates between actors. Four plans are generated, and two are chosen for further execution. During execution, 20 tool calls are allowed from the MCP server. Additionally, after five proof-checking calls, the critic model (Claude 3.7 Sonnet) is invoked and analyzes whether a deviation from the initial plan has occurred. We use Claude 3.5 Sonnet for the execution and re-planning, and Google Gemini Flash 2.0 for other tasks, due to the necessity of a big context. Results of the evaluation are shown in Table 2.
@@ 4.2 Agentic System
Table 2: Measuring the performance of different Rocq generation methods via CoqPilot
4.3 Ablation study
@@
Table 3: Ablation study of Multi-Agent Debate at planning stage
@@
Planning  Considering that software-verification tasks cannot be solved ad hoc, without explicit planning, we measure how removing the MAD layer and reverting to single-pass planning affects the proportion of successfully proved theorems. We run two versions of the agent: one generates plans via MAD, and the other produces a single plan in one LLM call without further refinement. Additionally, we include a configuration with the Planning stage entirely disabled, where the executor immediately attempts to construct a proof without any plan. This comparison shows that an agent without planning performs nearly identically, or slightly worse, than one guided by a poor single-pass plan, suggesting that a suboptimal plan does not provide advantage. In contrast, multi-step planning via MAD yields a consistent improvement across all groups, confirming the importance of structured plan refinement. An example of how MAD repairs a previously unsuccessful plan is presented in Appendix E.
RocqStar retrieval  To further justify the improvements of the agentic pipeline, we evaluate the agent with its retrieval component replaced by the baseline Jaccard index. The observed drop in performance compared to the full pipeline highlights the importance of the RocqStar retrieval approach, while the gain over the Claude 3.5 Sonnet model demonstrates the effectiveness of the proposed agentic design.
@@
5 Related Work
Many Rocq generation methods improve generation using Retrieval Augmentation. Most of those works solve the hint selection problem [3, 28], described in §˜2. Those approaches build proofs tactic by tactic, retrieving relevant lemmas or definitions to use in the next step. The problem of searching for existing proofs that could advance the generation is barely described in the literature. CoqPilot [15] and Rango [28] pack the context for the generator model with theorems most similar to the one we are solving. Our work proposes a novel method of doing premise selection and shows improvement over the baseline from previous works [15, 28].
In our Multi-Agentic system, we distribute responsibility over different agents. Differentiating between models that handle natural reasoning and those that handle coding is common practice in agentic systems. The work of Li et al. [17] proposes a similar task force split into Thinker, Solver, Critic, and Debug agents. Liang et al. [18] introduces a Multi-Agent debate framework, shows that this approach encourages divergent thinking, and demonstrates its usability in complex reasoning tasks. We show that planning is essential for the formal verification pipeline. Theorem-proving demands a clear, high-level picture of the proof before executing any code. Running a multi-agent debate at the planning stage ensures rigorous evaluation of different approaches before interacting with Rocq’s system. We produce several plans for further execution. In a manner, close to Islam et al. [11], we assign scores to plans and run them in order of score decrease. To our knowledge, there were close to no attempts to building Agentic systems for ITPs. Yang et al. [32] have shown an initial proof of concept of an agent for Lean; however, their agent lacks automaticity, the pipeline incorporates only minimal tooling, and does not possess an explicit planning stage.
As a user interface, we utilize CoqPilot to integrate into the common Rocq’s programmer pipeline. CoqPilot is a VSCode [^5] VSCode: https://code.visualstudio.com plugin, facilitating access to Rocq generation methods for end-users.
6 Conclusion
@@
Acknowledgments and Disclosure of Funding
We thank Ekaterina Verbitskaia, Ivan Kabashnyi, Maksim Rozenberg, and Pavel Guliaev for their valuable feedback on this work. We are also grateful to the JetBrains Koog team for providing a hands-on and well-designed framework for building AI agents. Finally, we thank the Dynamic Program Analysis Team (Egor Klimov, Nikita Dukin, and Saga Rut Sunnevudóttir) for their help in analyzing the agent’s execution traces.
Appendix A IMM Evaluation Dataset
The collected IMM-300 dataset from the CoqPilot [15] work includes only theorems with proofs of length no more than 20. For that reason the bucket with the most difficult theorems is labeled 9 — 20 tactics. This decision has been made, reflecting CoqPilot’s original focus on subgoals and shorter lemmas. Theorems of length no more than 20 tactics account for 83% of all proofs in the IMM project. As we take the same dataset, it possesses the same limitations. Therefore, we have not evaluated our solution on theorems, for which the reference proof contains more than 20 tactics. However, such theorems are quite rare.
The exact list of theorems used in each group could be found in the repository of the CoqPilot project: https://github.com/JetBrains-Research/coqpilot/blob/main/etc/docs/benchmark.
A common problem with testing pipelines that include general-purpose LLM providers, such as OpenAI, is data contamination. We are aware, that the model could have possibly seen the human-written proofs, as the IMM project has been publicly available since a while. However, firstly, the model sees neither the theorem name, for which it is generating the proof, nor the proof goal exactly as it was in the initial file. As we treat them as proof states, rather than theorems, an LLM receives it in an equivalent, but slightly modified way. Secondly, as many of our experiments have shown, various quality of premise selection drasticly changes the behavior of the model. That hints that the model is not able to memorize all theorems and proofs. Lastly, data contamination issue was one of the things we had in mind, while developing BigRocq. One could pass a Rocq project into BigRocq as input, and for each theorem retrieve the sub-state, that is achieved after $k$ steps. On an example of $k=2$, the following theorem:
Lemma eq_trans (A : Type) : forall (x y z : A), x = y -> y = z -> x = z.
Proof.
intros x y z Hxy Hyz.
rewrite Hxy. (* State: (A : Type) (x y z: A) (Hxy: x = y) (Hyz: y = z) : y = z *)
rewrite Hyz.
reflexivity.
Qed.
Could be automatically tranformed into the following one:
Proof.
rewrite Hyz.
reflexivity.
Qed.
Appendix B Encoder Training Dataset
@@
Due to the limited information we get from the Coq-LSP, our heuristic algorithm of transforming the proof into a tree breaks down. We cannot augment such theorems. The authors of CoqGym [31] also explicitly state that they do not handle theorems with goal selectors. They state that in their dataset, goal selectors occur in less than 1% data. The situation has changed since the work was published; the feature is now used more often but is still relatively rare. Goal selectors are an issue to be solved, and we are working on a solution by extracting some additional information from Rocq’s system through Coq-LSP.
@@
Dataset is split into training, validation, and test sets with proportions of 70%, 20%, and 10% respectively. Theorems from the same file do not appear in different sets. Parameters of building the dataset are listed in Table˜4b. Pair of statements is considered as negative, if the proof_distance between them is greater than 0.65, and positive if it is less than 0.3. If the distance is in range $[0.45,0.65]$, with probability of 30% it is also considered to be a negative pair (see hard negatives in §˜2.2).
Appendix C Encoder Details

### v3

@@ 3. Agentic System
Agent-based approaches are broadly used in code generation and repair tasks sweagent; autocoderover; repairagent. Despite a large number of autonomous and semi-autonomous coding agents, they are not widely used in formal proofs generation and are not tailored to the Rocq specifics. To address this, we have implemented a RocqStar agentic system.
@@
To allow interaction between the agent and Rocq’s system, we develop a REST API server that provides a set of tools that are useful during the execution. We apply our domain knowledge and construct these tools to bring an agent-driven proving process as close as possible to a human-driven one. Examples of allowed function calls include checking validity of proofs, retrieving the valid prefix of given proof, gathering additional information about available entities in the context, and interacting with the context via performing commands like Print ?a to identify the type of an argument or Search ?exp to search for defined terms by a pattern. Toolset is described in detail in Appendix C. Interaction with Rocq’s system is carried out through its language server, coq-lsp emiliocoqlsp. To conform with a commonly used Model Context Protocol (MCP) and allow seamless agent interaction with the environment through tools, we implement an MCP server that wraps the REST API server. Among the provided tools, the most important is the proof-checking tool. It not only verifies whether a proof is valid but, in case of an error, returns detailed diagnostic information: the error message, its exact location, the valid prefix preceding the error, and the remaining goals after that prefix. This functionality allows the agent to maintain awareness of the current proof state and leverage partial proof progress.
3.1. Agent Logic
@@
The input to the agent is presented as a target theorem without a proof and a file where it was declared, see box 1 of Fig. 2. Agent’s pipeline is logically split into two main stages: planning and execution. In the planning phase, multiple language models rigorously work out the strategy for the further implementation. During execution agents follow the plan aiming to generate the correct proof.
@@
Planning Stage  We employ the idea of multi-agent debates liang2023multiagenticdebate to generate a strategy for proving the given theorem. Specifically, two LLMs engage in a discussion: one proposes and defends an initial plan (pro LLM), while the other critiques it (con LLM); see box 2 in Fig. 2. After several debate rounds, the entire message history is passed to a judge LLM, which determines the winner and produces the final plan. By repeating this procedure, we generate $k$ candidate strategies. These are then evaluated by a plan scoring LLM that assigns each a numerical score (the higher, the better). Finally, the top-$l$ plans are selected and forwarded to the Execution Stage; see box 3 of Fig. 2.
@@
Execution Stage  For each of the selected plans, we run an executor agent that follows it step by step, invoking tools from the provided tool set — proof checker, context–inspection queries, search commands, and others, as atomic actions. Through these tool calls, the agent interacts with the environment via the MCP server. In addition to this iterative execution, we employ a reflection mechanism that monitors the progress of the proof and adjusts the strategy when necessary. We track how many consecutive erroneous proof attempts occur, and once this number exceeds a predefined threshold (set to five during evaluation), a critic model is called to assess the current proof state and identify deviations from the intended strategy. After that, we retrieve theorems along with their proofs, whose top-level goals are similar to the currently remaining goal, according to the cosine similarity between their RocqStar-ranker embeddings. We prompt the LLM to explain which tactic sequences could be helpful to finish our proof. We gather the generated criticism and send it to the replanner LLM to refine the current plan along with similar proofs and their analysis. The replanner is a separate language model that revises the plan based on the critic’s feedback and the retrieved examples. The whole message history is sent back to the executor agent. During the execution of each plan, $n$ tool calls are allowed. If valid proof is not found after $n$ tool calls, we denote the plan as failed. In this case, we ask a plan failure summarizer LLM to generate a short explanation of why the strategy execution failed and what happened during it. Then this summarized explanation is sent to the new execution stage with the next selected plan. This procedure is repeated until the correct proof is found or there are no more strategies to execute. For clarity, Figure Fig. 2 presents a simplified view of the system and does not explicitly depict all auxiliary LLM components used in the pipeline.
@@
Table 1. Model performance under different ablations across all evaluation sets
@@
Table 2. Different Rocq generation methods via CoqPilot
@@
Table 3. Ablation study of Multi-Agent Debate
4. Evaluation
To evaluate our approach, partially and as a whole, we use the CoqPilot benchmarking framework. We required a dataset with a large number of human-written theorems and proofs. To compare our solution to existing ones, we decided to re-use the dataset by coqpilot. It is limited to 300 theorems from the IMM project podkopaev2019bridging, which was suitable for us in terms of computational and financial costs. The theorems are partitioned into three groups, corresponding to the difficulty level. The length (in tactics) of the human-written reference proof of the theorem estimates its difficulty. The sizes of each group are chosen with respect to the initial distribution of proof lengths in the project. Final group sizes and length ranges of each group could be found in Table 3. The dataset is further restricted to theorems whose human-written reference proofs contain no more than 20 tactics, following the original setup of the CoqPilot benchmark. From now on, we will refer to the described dataset as the IMM-300 dataset. For smaller ablation studies we additionally prepared IMM-50, a 50-theorem subset of IMM, constructed with the same procedure. No theorems from the dataset were present in the training set of the RocqStar ranker embedding model. Moreover, the training set only contained partial theorem goals, i.e., intermediate proof goals arising during proof execution, rather than the original top-level theorem statements. The split of both datasets into groups, details, and limitations are described in Appendix D. Computational and financial resources used for experiments are described in Appendix E.
4.1. Retrieval Mechanism
We integrate our retrieval mechanism as a ranker into CoqPilot and evaluate it on the IMM-300 dataset with different models under the hood. To assess its performance, we compare our approach against two baselines: (i) an untrained embedder model (we use gte-modernbert-base, with codebert-base yielding comparable results), and (ii) a lexical similarity baseline based on the Jaccard index. In the latter, given a target theorem statement $s_{*}$ and a set of proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, it ranks the theorems in descending order of $J(s_{*},s_{i})$, where $J(s_{*},s_{i})$ is the Jaccard-similarity index. The statement is split into tokens by whitespaces, commas, etc. Jaccard-similarity index is semantically almost the same as the BM-25 metric and produces the same numerical results. For each theorem in the dataset, we take theorems within the same file, sort them using the ranker (Jaccard, ModernBert, or RocqStar, respectively), take the $k$ most relevant ones ($k$ is equal to 7 in our experiments) and send a request to the model to generate the completion. The chosen theorems are being sent as a few-shot prompt. Generation for each theorem is requested 12 times. If the Rocq’s system accepts any of the proofs, the theorem is considered solved. The target metric in our evaluation is the ratio of solved theorems. The evaluation results are presented in Table 1. The reported values denote mean success rates, and the $\pm$ intervals correspond to the standard deviation across three independent runs.
4.2. Agentic System
We evaluate our agentic system on the IMM-300 dataset with the goal of solving as many theorems as possible. We use the strongest available Rocq proof–generating models, as measured by pass@12 on the CoqPilot benchmark. In addition, following the common principle of using a critic stronger than the executor variationinverification, we employ a more capable model for the planning and evaluation stages. Specifically, for all parts of the planning stage, we use the Claude 3.5 Sonnet model and perform two rounds of debate between actors. Four plans are generated, and two are chosen for further execution. During execution, 20 tool calls are allowed from the MCP server. Additionally, after five proof-checking calls, the critic model (Claude 3.7 Sonnet) is invoked and analyzes whether a deviation from the initial plan has occurred. We use Claude 3.5 Sonnet for the execution and re-planning, and Google Gemini Flash 2.0 for other tasks, due to the necessity of a big context. Results of the evaluation are shown in Table 3.
4.3. Ablation study
@@
Planning  Considering that software-verification tasks cannot be solved ad hoc, without explicit planning, we measure how removing the MAD layer and reverting to single-pass planning affects the proportion of successfully proved theorems. We run two versions of the agent: one generates plans via MAD, and the other produces a single plan in one LLM call without further refinement. Additionally, we include a configuration with the Planning stage entirely disabled, where the executor immediately attempts to construct a proof without any plan. This comparison shows that an agent without planning performs nearly identically, or slightly worse, than one guided by a poor single-pass plan, suggesting that a suboptimal plan does not provide advantage. In contrast, multi-step planning via MAD yields a consistent improvement across all groups, confirming the importance of structured plan refinement. An example of how MAD repairs a previously unsuccessful plan is presented in Appendix F.
RocqStar retrieval  To further assess the contribution of retrieval to the agentic pipeline, we evaluate the agent with its retrieval component replaced by a Jaccard-based baseline. This substitution results in a small but consistent decrease in performance, indicating that the proposed retrieval mechanism provides a meaningful, though moderate, improvement in the agentic setting. While the agentic architecture mitigates some of the limitations of weaker premise selection, effective retrieval remains an important component that contributes to more stable and reliable proof generation.
@@
5. Related Work
Many Rocq generation methods improve generation using Retrieval Augmentation. Most of those works solve the hint selection problem knntactician; thompson2024rango, described in § 2. Those approaches build proofs tactic by tactic, retrieving relevant lemmas or definitions to use in the next step. The problem of retrieving existing proofs that can help advance generation is barely explored in the literature. CoqPilot coqpilot and Rango thompson2024rango address this problem by augmenting the generator’s context with previously proven theorems whose statements are similar to the target theorem being proved. Our work proposes a novel premise selection method and demonstrates improvements over the baselines used in prior work coqpilot; thompson2024rango.
In our multi-agentic system, we distribute responsibility across multiple agents, each responsible for a simpler subtask within the overall proof generation process. This decomposition, which separates high-level reasoning, planning, and execution, is a common design pattern in agentic systems. For example, li2024codetree propose a task force split into Thinker, Solver, Critic, and Debug agents, while liang2023multiagenticdebate introduce a multi-agent debate framework that encourages divergent reasoning in complex tasks. We show that explicitly separating planning from execution is essential for the formal verification pipeline. Theorem-proving demands a clear and high-level picture of the proof before executing any code. Running a multi-agent debate at the planning stage ensures rigorous evaluation of different approaches before interacting with Rocq’s system. We produce several plans for further execution. In a manner similar to islam2024mapcoder, we assign scores to plans and run them in the order of score decrease. To our knowledge, there have been no substantial attempts to build fully agentic systems for ITPs. The closest related effort is the early proof-of-concept by yang2023leandojo, which explores the use of agents for proof generation in Lean. While this work illustrates the potential of agent-based approaches in this domain, it relies on minimal tooling and does not form a fully agentic system with autonomous planning and iterative execution.
As a user interface, we utilize CoqPilot to integrate into the common Rocq’s programmer pipeline. CoqPilot is a VSCode [^5] VSCode: https://code.visualstudio.com plugin, facilitating access to Rocq generation methods for end-users. Currently, the RocqStar retrieval component is available in CoqPilot as a premise-ranking module. Integration of the full RocqStar agentic system as a proof generator is ongoing.
6. Conclusion
@@
{acks}
We thank Ekaterina Verbitskaia, Ivan Kabashnyi, Maksim Rozenberg, and Pavel Guliaev for their valuable comments on this work. We also thank the Dynamic Program Analysis Research team at JetBrains (Nikita Dukin, Saga Rut Sunnevudóttir, and Egor Klimov) for their help in building tool execution distributions described in Appendix G. Finally, we thank the Koog team at JetBrains for developing Koog, an open-source framework for building AI agents with an idiomatic, type-safe Kotlin DSL.
Appendix A Encoder Training Dataset
@@
Due to the limited information we get from the Coq-LSP, our heuristic algorithm of transforming the proof into a tree breaks down. We cannot augment such theorems. The authors of CoqGym coqgym also explicitly state that they do not handle theorems with goal selectors. They state that in their dataset, goal selectors occur in less than 1% data. The situation has changed since the work was published; the feature is now used more often but is still relatively rare. Goal selectors are an issue to be solved, and we are working on a solution by extracting some additional information from Rocq’s system through Coq-LSP.
@@
Dataset is split into training, validation, and test sets with proportions of 70%, 20%, and 10% respectively. Theorems from the same file do not appear in different sets. Parameters of building the dataset are listed in Table 4b. Pair of statements is considered as negative, if the proof_distance between them is greater than 0.65, and positive if it is less than 0.3. If the distance is in range $[0.45,0.65]$, with probability of 30% it is also considered to be a negative pair (see hard negatives in § 2.2).
Appendix B Encoder Details